OpenStudy (anonymous):

WILL MEDAL!!!! On which of the following lines does the point (7, 1) lie? a. y-5x=4 b. y=x-10 c. y=-4x+3 d. y=-x+8

OpenStudy (butterflydreamer):

hey :) A simple way to do this is just to substitute in your point (7, 1) into your equations. For example for the first equation : y - 5x = 4 You want to sub in x = 7 , y = 1 So, 1 - 4 ( 7) = 4 1 - 28 = 4 -27 = 4 is INCORRECT because they're not the same number :) Therefore, (7,1) does not satisfy y - 5x = 4 Basically you follow the same pattern and substitute x = 7 , y = 1 into your equations until you have the LHS = RHS ^_^